# ACC Saturator: Automatic Kernel Optimization for Directive-Based GPU Code

### v3

@@ IX Related Work
Numerous optimization techniques have been developed for GPU computing [57]. Rawat et al. [20] reorder stencil computation using heuristic algorithms to alleviate register pressure and increase SM occupancy. Software systolic arrays [58] assign global data to GPU threads, computing results by propagating and accumulating partial results through warp-level shuffles across overlapped blocks, improving data locality without shared memory usage. Hong et al. [59] utilize lightweight kernel emulation to provide auto-tuners with performance bottleneck feedback. In contrast, ACC Saturator adopts a straightforward rewriting mechanism, enabling its application to complex programs. Moreover, our method is complementary to other code optimizations.

### v4

@@ X Conclusion
Acknowledgement
The authors are funded by the EPEEC project from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 801051, the Ministerio de Ciencia e Innovación—Agencia Estatal de Investigación (PID2019-107255GB-C21/AEI/10.13039/501100011033), and Departament de Recerca i Universitats from the Generalitat de Catalunya as the Research Group AccMem (Code: 2021 SGR 00807). Antonio J. Peña was partially supported by the Ramón y Cajal fellowship RYC2020-030054-I funded by MCIN/AEI/10.13039/501100011033 and by “ESF Investing in your future”. The authors gratefully acknowledge the support of the NVIDIA Solutions Lab who provided us the remote access to their GPU environment. The authors would like to acknowledge the NVIDIA AI Technology Center (NVAITC) Europe for their valuable help. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525.